Blood is pumped from the heart at a rate of 9 L/min into the aorta (of radius 6.62 cm). Determine the speed of blood through the aorta.


Round your answer to 3 decimal places


1
Expert's answer
2021-04-29T10:45:39-0400

Answer

Flow rateQ is defined as

"Q=\frac{V}{t}=Av"

where V is the volume and t is the elapsed time.

So velocity is given by

"v=\frac{Q}{A}\\=\frac{9\times10^{-3}}{60\times3.14\times6.62\times6.62\times10^{-4}}"

"=0.0109m\/s"